A prism has a square base and a volume of 605dm3.If it is 12 dm high, what is the width of the base?
Answer by venugopalramana(3286)

A prism has a square base and a volume of 605dm3.If it is 12 dm high, what is the width of the base?
volume of prism=base area *ht.=base*base*ht
605=base^2*12
base = sqrt(605/12)=7.1 dm